OpenStudy (anonymous):

a building casts a 35-ft shadow and a flagpole casts a 10-ft shadow. If the flagpole is 22 ft tall, how tall is the building? Round to the nearest tenth. 77.0 ft 15.9 ft 37.0 ft 53.0

OpenStudy (anonymous):

proportionality

OpenStudy (anonymous):

use ratio theory 35/10 = x/22, so x = 77, which is the answer